1,4 butandiol is an organic alcohol that is made up of 53.3% carbon, 11.2% hydrogen, and 35.5% oxygen. It has a molar mass of 90.14 g/mol

a) what is the empirical formula based on the mass percentage

b) what is the molecular formula for 1,4 butandiol?


1
Expert's answer
2021-06-23T05:17:18-0400

a) C = (90.14 x 0.533) / 12 = 4

H = (90.14 x 0.112) / 1 = 10

O = (90.14 x 0.355) / 16 = 2

Empirical formula of 1,4 butandiol is C4H10O2


b) 1,4 butandiol - HOCH₂CH₂CH₂CH₂OH